OpenStudy (anonymous):

Find the axis of semetry and the vertex of the graph y=-x^2-10x+17

OpenStudy (anonymous):

\( y=-x^2-10x+17\) \( y=-(x^2+10x)+17\) \( y=-(x^2+10x +(10/2)^2)+17+(10/2)^2\) \( y=-(x^2+10x +(5)^2)+17+25\) \( y=-(x^2+5)^2+42\) \( y=-(x^2-(-5))^2+42\) The vertex is \(-5,42\)